A projectile (mass = 0.254 kg) is fired at and embeds itself in a target (mass = 1.76 kg). (The target is initially stationary). The target (with the projectile in it) flies off after being struck. What percentage of the projectile's incident kinetic energy does the target (with the projectile in it) carry off after being struck?
Answer #1

GIVEN

Let the Initial velocity of the projectile = u
Mass of projectile, m = 0.254 kg
Mass of target, M = 1.76 kg

Let the final velocity of the system be v.
Mass of the system = m + M = 2.014 kg

Apply the Principle Of Conservation of Momentum :
mu + 0 = (m + M)v
v = mu / (m + M)

Initial K.E. of the system = K.E. of projectile + K.E. of target = 1/2*mu^2 + 0 = 1/2*mu^2
Final K.E. of the system = 1/2*(m + M) v^2 = 1/2*m^2*u^2 / (m + M)

% of KE carried by the system = KE of system / KE of projectile * 100
= 100m / (m + M) = 100*0.254/(2.014)% =12.61%